OpenStudy (anonymous):

A law firm has eight male and six female lawyers. A committee of five, two male lawyers and three female lawyers, is to be chosen to represent the firm at a conference. How many committees are possible?

OpenStudy (jjuden):

the no. of ways of choosing 2 male from 8 is C(8,2)= 8*7/2 =28 and that of choosing 3 from 6 female lawyer is C(6,3)= 6*5*4/6= 20 hence reqd no. of ways is 28*20

OpenStudy (anonymous):

why dont you divide 6*5*4 by 3 instead of 6?

OpenStudy (jjuden):

try it

OpenStudy (anonymous):

i did, its not the right answer.

OpenStudy (jjuden):

ok try what i said

OpenStudy (anonymous):

i did and you get the right answer. but i am wondering where you got the "2" and "6" from when you divide